# Diffusion Model in Hyperspectral Image Processing and Analysis: A Review

### v1

@@ IV-B8 Data Recovery
Hyperspectral image (HSI) recovery aims to recover clean images from degraded observations [144], and the recovery of hyperspectral images plays a pivotal role in subsequent hyperspectral image applications. Hyperspectral data recovery faces challenges in effectively exploring the spatial nonlocal self-similarity and spectral low-rank properties inherent in HSI, while hyperspectral exists data scarcity. Existing model-based approaches have limitations in accurately modeling complex image features with manual a priori, and deep learning-based approaches have poor generalization capabilities, complex models, poor results, and parameter redundancy, Diffusion models, due to their impressive image recovery performance [145, 146, 147, 148, 149, 150, 151]. In particular, it has recently attracted attention in terms of noise robustness [92, 95, 140, 152, 115, 119]. Diffusion methods, however, are trained on large amounts of training data and perform very well within distributions, but may be quite susceptible to distributional bias.
@@
To address the above problem, Miao et al. proposed a self-supervised diffusion model for HSI recovery, the Denoising Diffusion Probabilistic Models for Spatial Spectroscopy (DDS2M) [138], which works by inferring the parameters of the proposed variational spatial spectral module (VS2M) during reverse diffusion, using only the degraded HSI without any additional training data. A loss function based on variational inference is customized to enable untrained spatial and spectral networks to learn a posteriori distributions that serve as transitions in the sampling chain to help reverse the diffusion process. Chen et al [56] use a dual window, iterative diffusion model, and reverse reconstruction to mitigate the impact of background estimation. Li et al. propose an untrained diffusion model with a pre-trained diffusion model (HIR-DIFF) as a A supervised HSI recovery framework [141], which can recover clean HIS from the product of two low-rank components, is proposed based on singular value decomposition (SVD) and rank-revealed QR (RRQR) decomposition, and a novel exponential noise schedule is proposed to speed up the recovery process. Li et al. propose a spectral latent diffusion enhancement module [136], which introduces a latent diffusion-enhanced rectangular transformer for HSI recovery. A multi-shape spatial rectangular self-attentive module is proposed, which leads to a tighter alignment with the desired HSI.
@@
Meanwhile, there are also data recovery problems like noise mixing characterized by signal correlated and signal uncorrelated components, visual perception is affected and discriminative image distortion. Si et al. proposed a new model of CASSIDiff to reconstruct CASSI metrics [153], which employs a feature fusion mechanism based on the Discrete Wavelet Transform (DWT) and a spatial-spectral attention mechanism, and has a high degree of reliability in the recovery task. reliability. Yi et al. proposed an efficient diffusion model (EDiffSR) [149] for RSI SR, developed an efficient activation network (EANet), and developed a practical conditional prior enhancement module (CPEM), which reduces the computational budget and retains more information. Meng et al. proposed a conditional diffusion model-based SRSISR method (FastDiffSR) [148], which reduces the number of sampling steps, employs residual images, and combines channel attention and spatial attention to obtain better visualization with significantly improved generation quality.
@@
V Challenges and Limitations of the Diffusion Model
V-A Computational complexity and Time cost of the model
Diffusion Models achieve data generation and denoising through multi-step iterative Markov chains, whose computational complexity increases exponentially with the data dimensions. Taking the denoising diffusion probabilistic model (DDPM) [19] as an example, hundreds of forward and reverse diffusion computations are required to generate a single hyperspectral image, especially when processing hyperspectral data above 200 bands, and the GPU memory occupancy can be more than 20 GB (e.g., 1024 × 1024 resolution scenes). Although the improved DDIM [35] reduces the number of sampling steps to less than 50 through non-Markov chains, the spatial-spectral three-dimensional nature of hyperspectral data (see Section II.B) still leads to a single inference time as long as several minutes, which makes it difficult to satisfy real-time demanding tasks, such as agricultural monitoring and disaster warning. Although the Latent Diffusion Model (LDM) [18] reduces the computational volume by compressing the latent space, the information compression in the spectral dimension may lose key features (e.g., mineral absorption peaks), which affects the classification and anomaly detection accuracy (see Table II).
V-B Impact of high dimensionality of hyperspectral data on model training
@@
V-C Adaptability issues to large-scale data
Real hyperspectral images often contain non-Gaussian noise (e.g., sensor thermal noise, atmospheric scattering noise), and most of the existing diffusion models are based on additive Gaussian noise assumptions (see Section III.A.1) The forward process design of the DDPM [19] is only applicable to isotropic Gaussian noise, and the denoising effect is significantly degraded for streak or PoissonGaussian hybrid noise (PSNR reduction of 8 15 dB). Although Score-SDE [21] extends the scope of noise modeling through stochastic differential equations, its inverse process relies on the accurate estimation of the score function, which is susceptible to error accumulation in bands where the signal-to-noise ratio (SNR) fluctuates drastically (e.g., near-infrared bands). For example, the Diffusion Models’ denoising performance degrades by about 20% compared to the conventional wavelet approach under extreme conditions where the noise standard deviation exceeds 30 dB. In addition, the mixed image element problem (see Section II.C) leads to a high degree of spatial coupling between the noise and the target signal, further exacerbating the modeling difficulty.
V-D Limitations of current models in practical applications
The black-box nature of diffusion models exposes them to credibility challenges in hyperspectral critical applications. For example, in anomaly detection tasks, the model may misclassify rare features as noise by overfitting the training data (literature [16]). Experiments have shown that LDM [18] based anomaly detectors have a 18% higher false alarm rate than the traditional RX algorithm (see Section II.C) in the detection of unknown mineral types. In addition, although the Latent Diffusion Model improves the computational efficiency, the physical meaning of the latent vectors is unclear, which makes it difficult for users to trace the generation logic of the detection results. In scenarios that require strict validation, such as military target identification, the interpretability disadvantage of diffusion models significantly constrains their practical application. Comparative studies have shown that traditional methods (e.g., SVM) [154] can provide a decision basis through feature importance analysis, whereas diffusion models can only output probability estimates and lack a transparent inference chain.
@@ V-D Limitations of current models in practical applications
VI Comparative and Experimental analysis
@@
TABLE II: Comparison of Diffusion Model for Hyperspectral Images with Traditional Deep Learning Models
@@
VI-A Experimental results and Performance assessment and Various metrics
VI-A1 Introduction to evaluation metrics and Datasets
@@
TABLE III: Summary of datasets used for the following experiments
@@
VI-A2 Experimental results
In this section, several typical methods from the above methods will be selected and compared to study the performance of detection, where traditional methods:GAN, VAE, autoregressive models, other deep learning methods, are compared with diffusion models. Since the experimental dataset of each algorithm is different, a common dataset is chosen for comparison as much as possible, in order to verify the anomaly detection performance of different algorithms on the same dataset in a more reasonable way. Where, the bolded indicates the algorithm’s performance optimal value for this dataset.
Comparison of classification experimental results, as shown in Figure 15, the classification effect of the first three methods is very poor, different classification color adulteration, the latter three methods overall classification effect is better, among them, the diffusion model-based method SpectralDiff classification effect is the best, followed by DMVL, their most boundary delineation is better versus the other methods, with less color adulteration, and SpectralDiff, despite the tiny classification there are some problems, but the overall effect is significantly better than other mainstream methods:
Figure 15: Comparison of different datasets on classification methods for hyperspectral images: (a) Pseudo-color image (b) Real labeled image (c) SF [155] (d) miniGCN [156] (e) SSFTT [157] (f) DMVL [158] (g) SSGRN [159] (h) SpectralDiff [108]
TABLE IV: Comparison of Metrics for Different Classification Methods on Indian Pines Dataset
@@ VI-A2 Experimental results
Table IV shows that the Spectral-Diff method is in the first place in Indian Pines dataset classification in terms of overall accuracy, average accuracy, and kappa coefficient, which is significantly improved over other methods and is the current optimal method.OA and AA are higher than the second place (SSGRN) by 0.8% and 0.34%, which indicates that the method is more robust in classification and more effective in capture features and reduce noise interference. However, it can also be seen that Spectral-Diff consumes much more time than other methods, which is its shortcoming.
TABLE V: Comparison of Metrics for Different Classification Methods on Pavia University Dataset
@@
As can be seen in Table V, the method has two metrics on the Pavia University dataset that are optimal 0.98% and 1.23% higher than the second (SSGRN), and the other is only 0.63% lower than the optimal, which is a relatively small gap, and it is still the optimal method among these methods, but the shortcoming of the use of samples is the long time consumed, which is more than twice as long as the other methods.
TABLE VI: Comparison of Metrics for Different Classification Methods on Salinas Dataset
@@
Table VI shows the classification on Salinas dataset and the same conclusion as the previous two tables, combined with Tables IV, V, and VI, the Spectral-Diff method is optimal in terms of performance compared to the other methods, but it is more time consuming and slower. SSGRN is optimal if balancing accuracy and efficiency, and miniGCN is optimal if low latency and high tolerance accuracy.
Comparison of target detection experimental results, as shown in Figure 16, the overall effect of these mainstream methods used in target detection is good, some methods exist undetected certain points or misdetection, among them, the diffusion model-based method Diff-Mosaic detection effect is the best, for the detection of the point of the edge of the border details of the best description, and the real labeling map is basically consistent, basically there is no missing parts:
Figure 16: Comparison of different datasets on target detection methods for hyperspectral images: (a) Original image (b) Real labeled image (c) ALC-Net [160] (d) ACM [161] (e) UIU-Net [162] (f) DNA-Net [163] (g) Diff-Mosaic [118]
TABLE VII: Comparison of Metrics for Different Target Detection Methods on SIRST Dataset
@@
Table VII shows that the Diff-Mosaic method has the optimal intersection ratio, detection probability, and false alarm rate for target detection on the SIRST dataset, which is 3.11% higher than the second place (DNA-Net), 1.89% higher than the second place (UIU-Net), and 87.8% lower than the second place (UIU-Net), respectively. Among them, the detection tends to 100% detection is very good with less error and is the best among these methods.
TABLE VIII: Comparison of Metrics for Different Target Detection Methods on NUDT-SIRST Dataset
@@
Table VIII shows the Diff-Mosaic in manual synthesis dataset NUDT-SIRST utilization, which is consistent with the metrics effect shown in Table 7, and moreover, it shows the superiority of this method among these methods with the best performance.
TABLE IX: Comparison of Parameters and Reasoning for Different Target Detection Methods
@@
Table IX shows that Diff-Mosai is less in the number of parameters, but not the best, while the inference time is the slowest among them, showing the shortcomings of the method in applications, 57.3% higher than the fastest time (ACM). Meanwhile, combining Tables VII, VIII and IX, again Diff-Mosaic is the best if pursuing the best performance, DNA-Net is the best if balancing performance and efficiency, and ACM is the best if high tolerance and low latency.
Comparison of the results of the anomaly detection experiments, as shown in Figure 17, compared with the traditional mainstream methods, the diffusion model-based methods BSDM, DWSDiff in the anomaly detection of the best detection effect, high detection rate, low false detection rate, and the real labeling map is basically consistent with the other mainstream detection methods to detect the detection of the detection effect is poor, the detection rate is low (the existence of the detection of the anomalies do not exist), the false detection rate is high:
Figure 17: Comparison of different datasets on hyperspectral image anomaly detection methods: (a) Pseudo-color map (b) Real label map (c) GAED [164] (d) RGAE [165] (e) Auto-AD [166] (f) DFAN-HAD [167] (g) DirectNet [168] (h) BSDM [133] (i) TGFA [169] (j)DWSDiff [56]
TABLE X: Comparison of metrics for each anomaly detection methods
@@
TABLE XI: Comparison of metrics for each anomaly detection methods
@@
TABLE XII: Comparison of metrics for each anomaly detection methods
@@
As can be seen in Table X, this table shows the accuracy of the anomaly detection methods, in which BSDM and DWSDiff are based on diffusion model, DWSDiff is in the first place in all other items except for the integrated detection performance (${mAUC}_{TD}$) which is second only to DFAN, while the data of BSDM is not outstanding (third in the integrated detection performance), but it shows the effect is not is not bad.
TABLE XIII: Comparison of Reasoning Time for Each Anomaly Detection Method
@@
Table XI shows that both diffusion models have longer and slower inference times for each medium dataset scenario, with average times 99.8% higher (BSDM) and 99.7% higher (DWSDiff) than the shortest time (Auto-Net), which is hundreds of times higher than the optimal time, and is very slow in comparison. Again seeking performance DWSDiff is optimal and BSDM is a good choice.
Comparison of the experimental results of image recovery, as shown in Fig. 18, the overall effect of image recovery are relatively modest, but the methods based on the diffusion model (DDPM, EDiffSR, FastDiffSR) are better handled in terms of details, and the degree of recovery is closer to the high-resolution map:
Figure 18: Comparison of different datasets on hyperspectral image recovery methods: (a) High resolution map (b) HESNet [170] (c) TransENet [171] (d) NDSRGAN [172] (e) HAT [173] (f) DDPM [19] (g) EDiffSR [149] (h) FastDiffSR [148]
TABLE XIV: Comparison of Metrics for Different Image Recovery Methods on Potsdam Dataset
@@
As can be seen in Table XII, the three methods based on diffusion model (DDPM, EDiffSR, FastDiffSR) show unique advantages in key indicators, among which, the FsatDiffSR method has the best overall performance in terms of PSNR and FID, which is far superior to the traditional methods (compared to the second: (HAT) 0.43%, (DDPM) 6.07%, and (DDPM) 13.5% difference), indicating that its recovered image reaches a higher level in both detail fidelity and distributional consistency, but in terms of the number of parameters, the diffusion model-based methods are all higher than HESNet: 79.4%, 79.7%, 76.6%, respectively, thus affecting the training time too long.
TABLE XV: Comparison of Metrics for Different Image Recovery Methods on Toronto Dataset
@@
Similar to Table XII, it can be seen from Table XIII that the three methods based on the diffusion model have a good overall performance, especially FastDiffSR has the best overall performance, and combining the two tables, the three diffusion methods have a higher quality of generation and more semantic information retention than the traditional methods. However, the three methods still have the problem of larger number of parameters in comparison.
TABLE XVI: Comparison of Measurement metrics for Different Image Recovery Methods
@@
The entropy and generation quality metrics in Table XIV show that the three diffusion methods retain more semantic information and have higher generation and recovery quality, corresponding to Tables XII and XIII.Also combining the three tables, overall FastDiffSR has a balanced performance in terms of recovery accuracy and efficiency, DDPM is more advantageous in terms of perceptual quality, and EDiffSR can be used as a practical choice for moderate performance.
Comparison of noise suppression experimental results, as shown in Figure 19, with the noise map as well as the real noiseless map compared to the overall performance of the Diff-Unmix method has the best effect, followed by DDRM, and finally DDS2M are based on the diffusion model method, they are more successful noise suppression; in contrast to the traditional mainstream method effect is worse, to LLxRGTV in which the best:
Figure 19: Comparison of different datasets on hyperspectral noise suppression methods: (a) Noise map (b) True noise-free map (c) LLxRGTV [174] (d) FGSLR [175] (e) LRTDCTV [176] (f) DIP [177] (g) LLnRPnP [178] (h) HLRTF [110] (i) DDRM [179] (j)DDS2M [138] (k)Diff-Unmix [142]
TABLE XVII: Comparison of Metrics for Different Noise Suppression Methods on KAIST Dataset
Methods Evaluation metrics(KAIST Gaussian noise: N(0,0,3))
@@
TABLE XVIII: Comparison of Metrics for Different Noise Suppression Methods on KAIST Dataset
Methods Evaluation metrics(KAIST Gaussian noise: N(0,0,2))
@@
The last three methods (DDRM/DDS2M/Diff-Unmix) in the experiment are based on the diffusion model, as can be seen from Tables XV and XVI, the first choice of the KAIST two different noise situations in the dataset, these three methods in the noise suppression performance have their own characteristics, and the Diff-Unmix in the key indexes is the most outstanding performance, which PSRN, SSIM, and FSIM are the highest values, compared with other methods to show its significant advantages in noise suppression, structure preservation and spectral fidelity, but the time processing of the method under the diffusion model is slightly higher (HLRTF is the fastest), and the overall efficiency is still in the acceptable range.
TABLE XIX: Comparison of Metrics for Different Noise Suppression Methods on CAVE Dataset
@@
Table XVII shows that the methods under the diffusion model are overall better than the other methods, of which Diff-Unmix remains the most prominent, demonstrating that the diffusion model method is overall superior to the other methods, with the best performance and the noise suppression presenting the optimal effect.
TABLE XX: Comparison of Metrics for Different Noise Suppression Methods on CAVE-Toy Dataset

### v2

@@ 4.2.8 Data Recovery
Hyperspectral Image (HSI) recovery aims to reconstruct clean images from degraded observations [144], playing a pivotal role in subsequent HSI applications. This recovery process faces significant hurdles in effectively leveraging the intrinsic spatial nonlocal self-similarity and spectral low-rank properties of HSIs, further complicated by existing data scarcity. Traditional model-based approaches struggle to accurately model complex image features with manual prior assumptions, while deep learning methods often suffer from poor generalization, model complexity, suboptimal results, and parameter redundancy. However, diffusion models have recently garnered considerable attention for their impressive image recovery performance [145, 146, 147, 148, 149, 150, 151], particularly for their robustness to noise [92, 95, 140, 152, 115, 119]. It’s worth noting, though, that diffusion methods are typically trained on large datasets and perform exceptionally well within those distributions, but they can be quite susceptible to distributional bias.
@@
To tackle these issues, Miao et al. introduced DDS2M (Denoising Diffusion Probabilistic Models for Spatial Spectroscopy) [138], a self-supervised diffusion model for HSI recovery. This model operates by inferring the parameters of its proposed variational spatial spectral module (VS2M) during reverse diffusion, using only the degraded HSI without needing any additional training data. A custom loss function based on variational inference enables untrained spatial and spectral networks to learn a posteriori distributions, which serve as transitions in the sampling chain to aid the reverse diffusion process. Chen et al. [56] employ a dual window, iterative diffusion model, and reverse reconstruction to mitigate the impact of background estimation. Li et al. proposed HIR-DIFF (HSI Recovery via Diffusion Model with Untrained Prior) [141], a supervised HSI recovery framework. This framework, based on singular value decomposition (SVD) and rank-revealed QR (RRQR) decomposition, recovers clean HSIs from the product of two low-rank components. They also introduced a novel exponential noise schedule to accelerate the recovery process. Furthermore, Li et al. presented a spectral latent diffusion enhancement module [136], which incorporates a latent diffusion-enhanced rectangular transformer for HSI recovery. Their proposed multi-shape spatial rectangular self-attentive module leads to a tighter alignment with the desired HSI.
@@
Meanwhile, in the realm of data recovery, problems often arise from noise mixing, characterized by both signal-correlated and signal-uncorrelated components, which can affect visual perception and cause discriminative image distortion. To address this, Si et al. proposed CASSIDiff [153] for reconstructing CASSI metrics. This model employs a feature fusion mechanism based on the Discrete Wavelet Transform (DWT) and a spatial-spectral attention mechanism, demonstrating high reliability in recovery tasks. Yi et al. introduced an efficient diffusion model (EDiffSR) [149] for RSI (Remote Sensing Image) Super-Resolution. They developed an efficient activation network (EANet) and a practical conditional prior enhancement module (CPEM), which collectively reduce the computational budget while retaining more information. Meng et al. proposed FastDiffSR [148], a conditional diffusion model-based SRSISR (Single-Image Super-Resolution) method. FastDiffSR significantly improves generation quality and visual clarity by reducing the number of sampling steps, utilizing residual images, and combining channel and spatial attention mechanisms.
@@
5 Challenges and Limitations of the Diffusion Model
5.1 Computational complexity and Time cost of the model
Diffusion Models achieve data generation and denoising through multi-step iterative Markov chains, whose computational complexity increases exponentially with the data dimensions. Taking the denoising diffusion probabilistic model (DDPM) [19] as an example, hundreds of forward and reverse diffusion computations are required to generate a single hyperspectral image, especially when processing hyperspectral data above 200 bands, and the GPU memory occupancy can be more than 20 GB (e.g., 1024 × 1024 resolution scenes). Although the improved DDIM [35] reduces the number of sampling steps to less than 50 through non-Markov chains, the spatial-spectral three-dimensional nature of hyperspectral data (see Section 2.2) still leads to a single inference time as long as several minutes, which makes it difficult to satisfy real-time demanding tasks, such as agricultural monitoring and disaster warning. Although the Latent Diffusion Model (LDM) [18] reduces the computational volume by compressing the latent space, the information compression in the spectral dimension may lose key features (e.g., mineral absorption peaks), which affects the classification and anomaly detection accuracy (see Table 2).
5.2 Impact of high dimensionality of hyperspectral data on model training
@@
5.3 Adaptability issues to large-scale data
Real hyperspectral images often contain non-Gaussian noise (e.g., sensor thermal noise, atmospheric scattering noise), and most of the existing diffusion models are based on additive Gaussian noise assumptions (see Section 3.1.1) The forward process design of the DDPM [19] is only applicable to isotropic Gaussian noise, and the denoising effect is significantly degraded for streak or PoissonGaussian hybrid noise (PSNR reduction of 8 15 dB). Although Score-SDE [21] extends the scope of noise modeling through stochastic differential equations, its inverse process relies on the accurate estimation of the score function, which is susceptible to error accumulation in bands where the signal-to-noise ratio (SNR) fluctuates drastically (e.g., near-infrared bands). For example, the Diffusion Models’ denoising performance degrades by about 20% compared to the conventional wavelet approach under extreme conditions where the noise standard deviation exceeds 30 dB. In addition, the mixed image element problem (see Section 2.3) leads to a high degree of spatial coupling between the noise and the target signal, further exacerbating the modeling difficulty.
5.4 Limitations of current models in practical applications
The black-box nature of diffusion models exposes them to credibility challenges in hyperspectral critical applications. For example, in anomaly detection tasks, the model may misclassify rare features as noise by overfitting the training data (literature [16]). Experiments have shown that LDM [18] based anomaly detectors have a 18% higher false alarm rate than the traditional RX algorithm (see Section 2.3) in the detection of unknown mineral types. In addition, although the Latent Diffusion Model improves the computational efficiency, the physical meaning of the latent vectors is unclear, which makes it difficult for users to trace the generation logic of the detection results. In scenarios that require strict validation, such as military target identification, the interpretability disadvantage of diffusion models significantly constrains their practical application. Comparative studies have shown that traditional methods (e.g., SVM) [154] can provide a decision basis through feature importance analysis, whereas diffusion models can only output probability estimates and lack a transparent inference chain.
@@ 5.4 Limitations of current models in practical applications
6 Comparative and Experimental analysis
@@
Table 2: Comparison of diffusion and traditional deep learning models for HSI
@@
6.1 Experimental results and Performance assessment and Various metrics
6.1.1 Introduction to evaluation metrics and Datasets
@@
Table 3: Description of benchmark datasets used for evaluation
@@
6.1.2 Experimental results
In this section, a selection of typical methods from those previously discussed will be chosen for a comparative study of their detection performance. This comparison will encompass traditional methods such as GANs, VAEs, autoregressive models, other deep learning approaches, and diffusion models. To ensure a more reasonable assessment of anomaly detection performance across different algorithms, a common dataset will be used for comparison whenever possible. In the results, the bolded values will indicate the optimal performance achieved by an algorithm on that specific dataset.
As illustrated in Figure 15, the classification performance of the first three methods is notably poor, exhibiting significant color adulteration across different classes. In contrast, the latter three methods demonstrate superior overall classification effects. Among these, the diffusion model-based method, SpectralDiff, achieves the best classification results, closely followed by DMVL. Both SpectralDiff and DMVL show improved boundary delineation compared to other methods, with minimal color adulteration. Although SpectralDiff has some minor classification issues in tiny areas, its overall performance is significantly better than other mainstream approaches.
Figure 15: Classification results on multiple datasets using different methods: (a) Pseudo-color image (b) Real labeled image (c) SF [155] (d) miniGCN [156] (e) SSFTT [157] (f) DMVL [158] (g) SSGRN [159] (h) SpectralDiff [108]
Table 4: Comparison of Metrics for Different Classification Methods on Indian Pines Dataset
@@ 6.1.2 Experimental results
Table 4 reveals that the Spectral-Diff method achieves the top position in the Indian Pines dataset classification across overall accuracy (OA), average accuracy (AA), and kappa coefficient. Its performance shows a significant improvement over other methods, marking it as the current optimal approach. Notably, its OA and AA are 0.8% and 0.34% higher than the second-place method (SSGRN), indicating superior robustness in classification, more effective feature capture, and reduced noise interference. However, a drawback of Spectral-Diff is its considerably higher time consumption compared to other methods.
Table 5: Comparison of Metrics for Different Classification Methods on Pavia University Dataset
@@
As Table 5 shows, this method achieves optimal results on the Pavia University dataset for two metrics, surpassing the second-place method (SSGRN) by 0.98% and 1.23%. While another metric is only 0.63% lower than the optimum, which is a relatively small difference, it remains the best among these methods. However, a significant drawback is its time consumption, which is more than double that of other approaches.
Table 6: Comparison of Metrics for Different Classification Methods on Salinas Dataset
@@
Table 6 presents the classification results on the Salinas dataset, leading to the same conclusion as Tables 4 and 5. Across all three datasets (Indian Pines, Pavia University, and Salinas), the Spectral-Diff method consistently demonstrates optimal performance when compared to other approaches. However, this superior accuracy comes at the cost of higher computational time and slower processing speeds. If the goal is to balance accuracy with efficiency, SSGRN emerges as the optimal choice. Conversely, for applications demanding low latency and high tolerance for accuracy, miniGCN is the most suitable option.
As seen in Figure 16, the overall performance of these mainstream target detection methods is good. However, some methods exhibit instances of undetected points or false positives. Among them, the diffusion model-based method, Diff-Mosaic, demonstrates the best detection effect, providing the most accurate and detailed boundary descriptions for detected points. It is largely consistent with the true labeling map, with virtually no missing components.
Figure 16: Target detection results on various datasets with competing methods: (a) Original image (b) Real labeled image (c) ALC-Net [160] (d) ACM [161] (e) UIU-Net [162] (f) DNA-Net [163] (g) Diff-Mosaic [118]
Table 7: Comparison of Metrics for Different Target Detection Methods on SIRST Dataset
@@
Table 7 demonstrates that the Diff-Mosaic method achieves optimal performance for target detection on the SIRST dataset. It surpasses the second-place methods (DNA-Net and UIU-Net) with a 3.11% higher intersection ratio and a 1.89% higher detection probability. Furthermore, its false alarm rate is 87.8% lower than UIU-Net. This indicates that Diff-Mosaic’s detection rate approaches 100% with minimal errors, making it the superior method among those compared.
Table 8: Comparison of Metrics for Different Target Detection Methods on NUDT-SIRST Dataset
@@
Table 8 further corroborates the findings from Table 7, showing that Diff-Mosaic consistently delivers optimal performance on the manually synthesized NUDT-SIRST dataset. This reinforces the method’s superior effectiveness among the compared approaches.
Table 9: Comparison of Parameters and Reasoning for Different Target Detection Methods
@@
Table 9 indicates that while Diff-Mosaic has a comparatively lower number of parameters, it isn’t the most parameter-efficient method. More significantly, its inference time is the slowest among those compared, being 57.3% higher than the fastest method (ACM). This highlights a key limitation of Diff-Mosaic in practical applications where speed is critical. When considering the collective results from Tables 7, 8, and 9, the optimal choice for target detection depends on specific priorities. If the goal is to achieve the absolute best performance, Diff-Mosaic stands out. However, if balancing strong performance with efficiency is paramount, DNA-Net proves to be the most suitable option. Conversely, for applications demanding high tolerance for minor accuracy variations and low latency, ACM is the ideal choice.
As presented in Figure 17, when comparing the results of anomaly detection experiments, the diffusion model-based methods, BSDM and DWSDiff, demonstrate superior detection performance compared to traditional mainstream approaches. They achieve a high detection rate and a low false detection rate, largely aligning with the true labeling map. In contrast, other mainstream detection methods exhibit poorer detection effects, characterized by lower detection rates (missing existing anomalies) and higher false detection rates.
Figure 17: Comparative anomaly detection performance across hyperspectral datasets: (a) Pseudo-color map (b) Real label map (c) GAED [164] (d) RGAE [165] (e) Auto-AD [166] (f) DFAN-HAD [167] (g) DirectNet [168] (h) BSDM [133] (i) TGFA [169] (j)DWSDiff [56]
Table 10: Comparison of metrics for each anomaly detection methods
@@
Table 11: Comparison of metrics for each anomaly detection methods
@@
Table 12: Comparison of metrics for each anomaly detection methods
@@
As Table 10, 11, 12 illustrates, which presents the accuracy of anomaly detection methods, BSDM and DWSDiff are diffusion model-based approaches. Among these, DWSDiff ranks first in all metrics except for integrated detection performance (${mAUC}_{TD}$), where it is second only to DFAN. While BSDM’s data isn’t as outstanding (ranking third in integrated detection performance), its effectiveness is still commendable.
Table 13: Comparison of Reasoning Time for Each Anomaly Detection Method
@@
Table 13 indicates that both diffusion models, BSDM and DWSDiff, exhibit significantly longer and slower inference times across various medium dataset scenarios. Their average times are 99.8% (BSDM) and 99.7% (DWSDiff) higher than the shortest time achieved by Auto-Net, translating to hundreds of times slower than the optimal. This highlights a considerable challenge in terms of computational efficiency. Nonetheless, if the primary goal is to maximize anomaly detection performance, DWSDiff remains the optimal choice, with BSDM also being a strong contender.
As depicted in Figure 18, the overall image recovery effects are relatively modest across all methods. However, diffusion model-based approaches (DDPM, EDiffSR, FastDiffSR) demonstrate superior handling of details, resulting in a recovery quality that is notably closer to the high-resolution ground truth.
Figure 18: Visual comparison of hyperspectral image recovery methods: (a) High resolution map (b) HESNet [170] (c) TransENet [171] (d) NDSRGAN [172] (e) HAT [173] (f) DDPM [19] (g) EDiffSR [149] (h) FastDiffSR [148]
Table 14: Comparison of Metrics for Different Image Recovery Methods on Potsdam Dataset
@@
As evidenced in Table 14, the three methods based on diffusion models—DDPM, EDiffSR, and FastDiffSR—demonstrate unique advantages in key image recovery indicators. Notably, FastDiffSR exhibits the best overall performance in terms of PSNR and FID, significantly outperforming traditional methods. Compared to the second-best methods (HAT for PSNR and DDPM for FID), FastDiffSR shows substantial improvements: 0.43% for PSNR over HAT, and 6.07% and 13.5% for FID over DDPM. This indicates that FastDiffSR recovers images with superior detail fidelity and distributional consistency. However, a drawback for all diffusion model-based methods is their higher number of parameters compared to HESNet, specifically 79.4% higher for DDPM, 79.7% for EDiffSR, and 76.6% for FastDiffSR. This increased parameter count contributes to longer training times.
Table 15: Comparison of Metrics for Different Image Recovery Methods on Toronto Dataset
@@
Similar to Table 14, Table 15 also demonstrates that the three methods based on the diffusion model exhibit strong overall performance, with FastDiffSR again standing out as the best. When combining insights from both tables, it’s clear that these three diffusion methods produce higher quality generations and retain more semantic information compared to traditional approaches. However, a consistent issue across all three diffusion methods remains their relatively larger number of parameters.
Table 16: Comparison of Measurement metrics for Different Image Recovery Methods
@@
Table 16, which presents entropy and generation quality metrics, reinforces the findings from Tables 14 and 15. It demonstrates that the three diffusion methods—FastDiffSR, DDPM, and EDiffSR—retain more semantic information and achieve higher generation and recovery quality. Overall, FastDiffSR offers a balanced performance in terms of both recovery accuracy and efficiency. DDPM holds an advantage in perceptual quality, while EDiffSR can be considered a practical choice for moderate performance needs.
As illustrated in Figure 19, when comparing the noise map and the true noiseless map, the Diff-Unmix method demonstrates the best overall performance in noise suppression, followed by DDRM and then DDS2M. All of these are diffusion model-based approaches, and they achieve notably successful noise suppression. In contrast, traditional mainstream methods exhibit poorer effects, with LLxRGTV being the best among them.
Figure 19: Performance comparison of denoising methods on hyperspectral images: (a) Noise map (b) True noise-free map (c) LLxRGTV [174] (d) FGSLR [175] (e) LRTDCTV [176] (f) DIP [177] (g) LLnRPnP [178] (h) HLRTF [110] (i) DDRM [179] (j)DDS2M [138] (k)Diff-Unmix [142]
Table 17: Comparison of Metrics for Different Noise Suppression Methods on KAIST Dataset
Methods Evaluation metrics(KAIST Noise: N(0,0,3))
@@
Table 18: Comparison of Metrics for Different Noise Suppression Methods on KAIST Dataset
Methods Evaluation metrics(KAIST Noise: N(0,0,2))
@@
In the experiments detailed in Tables 17 and 18, the last three methods—DDRM, DDS2M, and Diff-Unmix—are all based on the diffusion model. When evaluated on the KAIST dataset under two different noise conditions, these three methods demonstrate distinct strengths in noise suppression. Diff-Unmix consistently shows the most outstanding performance across key indices, achieving the highest PSNR, SSIM, and FSIM values. This highlights its significant advantages in noise suppression, structure preservation, and spectral fidelity compared to other methods. While the processing time for diffusion model-based methods like Diff-Unmix is slightly higher (with HLRTF being the fastest), their overall efficiency remains within an acceptable range.
Table 19: Comparison of Metrics for Different Noise Suppression Methods on CAVE Dataset
@@
Table 19 indicates that the diffusion model-based methods generally outperform other approaches, with Diff-Unmix remaining the most prominent. This consistently demonstrates the overall superiority of diffusion models in noise suppression, delivering optimal performance and the best noise suppression effects.
Table 20: Comparison of Metrics for Different Noise Suppression Methods on CAVE-Toy Dataset